A planet of mass m describes an elliptical orbit around the sun. The areal velocity of the planet is proportional to 

Grade:12

1 Answers

Arun
5 years ago
Kepler's law of Areas states that a line that connects a planet to the sun sweeps out equal areas in equal times.
mathematically,
dA/dt =L/2m (as areal velocity=L/2m = dA/dt]
 
L = mvr sin theta 
 
dA/dt = L/ 2m = vr sin theta / 2
 
hence independent of mass